# Implementing feature binding through dendritic networks of a single neuron

### v1

@@ II-C Data analysis
We simulated the PC morphology model in NEURON 8.4. The time step of all simulation experiments was set at 0.025ms. Data collected after simulations was saved for further analysis. In order to avoid the influence of ionic current on the nonlinear synaptic integration ability of dendrites, there is no ion channel distribution in all compartments of PC when calculating dendritic nonlinearity. To study dendritic nonlinearity, we record observed EPSPs ($\rm EPSP_{O}$) and the EPSPs produced by individual synapses are added together to obtain the expected EPSP ($\rm EPSP_{E}$). The sublinearity index was measured as the average sublinear summation $\rm(1-(EPSP_{O}/EPSP_{E}))*100$ with the last two points as in the dash box in Figure 1. To investigate the temporal dependence of dendritic integration, we systematically varied the inter-spike intervals (ISI) between synaptic inputs. Then, the amplitude of all EPSPs induced by the stimuli was averaged to obtain the EPSP value elicited by the stimulus. From this, the observed values and expected values were obtained, and the sublinearity index was calculated.
@@
$\rm F(EPSP_{E})=\frac{EPSP_{O}{max}}{1+(EPSP_{E}{50}/EPSP_{E})^{n}}+EPSP_{O}{0}$ (10)
where $\rm n$ is the exponent factor, $\rm EPSP_{O}{0}$ the observed EPSP offset, and $\rm EPSP_{O}{max}$ the maximum observed EPSP. $\rm EPSP_{E}{50}$ is the value of expected EPSP at which observed EPSP reaches half maximum. To characterize EPSP in detail, we introduced three measures, the peak, rise time, and width of EPSP. Peak is the maximum peak value of EPSP, and width as the distance between the points where the half-amplitude reaches before and after the peak.
In order to further study the effect of dendritic nonlinearity on PC firing. The dendrites with less than $20\%$ sublinearity are regarded as main dendrites, and those with more than $20\%$ are regarded as spiny dendrites, and then the spiny dendrites are divided into 24 branches, and the parent node of each branch is the main dendrite.
Feature binding refers to the brain’s ability to integrate various attributes (e.g., color, shape, motion) into a coherent percept. This process involves complex, nonlinear interactions within neuronal dendrites. When there are multiple synaptic inputs $\rm I_{1},I_{2},\ldots,I_{n}$ The nonlinear response $\rm R$ can be described by a function $\rm f$:
$\begin{split}R=f(I_{1}+I_{2}+\ldots+I_{n})\end{split}$ (11)
When studying PC computing performance, each branch is considered a computing unit. When the synapses distributed on the branch are activated, the input $\rm I_{n}$ is considered to be 1, and when it is not activated, the input $\rm I_{n}$ is considered to be 0. When studying Boolean functions, choose any 2 from 24 branches as $\rm X_{1}$ and $\rm X_{2}$. when studying the feature binding problem, 4 of the 24 branches are selected as $\rm X_{1}$, $\rm X_{2}$, $\rm X_{3}$, and $\rm X_{4}$.
@@ III-A PC dendrites show sublinear integration
Figure 1: PC individual dendrites show sublinear integration to the somatic response. (A) (Left) 5 synapses are distributed on thin (blue) and thick (red) dendrites. (Right) Single EPSPs evoked by single pulse synaptic stimulation; Observed EPSPs evoked with increasing synapses; Expected EPSPs as the algebraic sum of single EPSPs. (B) Sublinear relationship of the expected and observed EPSPs from (A). Dotted line represents linear function. The sublinearity index, 49% (thin) and 15% (thick) was measured as the average sublinear summation in the dashed box. Solid line indicates fits by Hill function. (C) The distributed sublinearity index for each dendrite. Sublinearity as a function of dendritic diameter and distance from soma.
@@
Figure 2: Temporal dependence of sublinear dendritic integration. (A) (Left) Synapses are distributed on thin (blue) and thick (red) dendrites. (Right) Single EPSPs evoked by six-pulse synaptic stimulation; Observed EPSC with increasing synapses; Expected EPSP as algebraic sum of single EPSPs. The pulse interval is 5 ms. (B) Sublinear relationship of normalized expected and observed EPSPs at different time intervals of 1, 5, and 10 ms. The dotted line represents linear function. The solid line indicates fits by Hill function. (C) The distributed sublinearity index for each dendrite when the pulse interval is 1, 5, and 10 ms. Sublinearity as a function of dendritic diameters.
@@
To examine whether the large dendritic depolarization could decrease the synaptic current driving force and introduce a nonlinearity that would curtail linear summation of somatic EPSPs (excitatory postsynaptic potentials) within the same PC dendrite, we studied subthreshold input-output relationships by comparing the algebraic sum of individual EPSP from five activated synapses along different yet nearby spatial locations along one dendrite, with the compound EPSP in response to simultaneous activation of multiple 2 to 5 locations (Figure 1A). The compound EPSP observed ($\rm{EPSP_{O}}$) was systematically smaller than the algebraic sum of its corresponding individual EPSP expected ($\rm{EPSP_{E}}$), especially on thin dendrites. The input-output relationship of dendritic summation using these five synapses shows sublinear trends that can be fitted by the Hill function. The sublinearity index (SI) can be calculated (the average of $\rm{(1-(EPSP_{O}/EPSP_{E}))*100}$ using the last two points; dashed box in Figure 1B) for each dendrite. The SI of the thin dendrite (diameter $0.29\mu m$, 49%) is stronger than the thick dendrite (diameter $1.75\mu m$, 15%) as in Figure 1B. We then asked how dendritic properties regulate the SI (Figure 1C). The SI of 739 dendrites was distributed over the entire dendritic field, which shows apical dendrites exhibit stronger SIs, but trunk dendrites are weaker. SI decreases with larger diameter dendrites while increasing with the dendritic distance from the soma. Such observations are inherent in the observed and expected EPSPs (Figure S1). The peak, rise time, and width of EPSP are also location dependent (Figure S2).
@@ III-B Strong sublinearity prefers a global scatter input strategy
Figure 3: Scattered synaptic inputs on strong sublinear dendrites trigger larger somatic responses. (A) 24 branches (color-coded) with strong sublinearity are segmented in the entire morphology. Each branch groups a set of single dendrites with sublinearity index ¿20%. Black trunk represents the main dendrites (dendritic sublinearity ¡20%). (B) Somatic EPSP in response to 50 synaptic inputs in a branch. Individual gray traces represent the response to a cluster input (50 synapses are distributed closely) on a single dendrite. The color trace shows synapses scattered throughout the branch. (C) Illustration of the somatic spiking threshold $\theta$ with varying degrees (solid lines) of dendritic sublinearity. Strong sublinearity makes soma voltage more difficult to reach the spiking threshold. (D) Input-output (IO) relationship curves changing over the sum of dendritic branches. (Left) IO curves in branch 1 (br1 in (B)). Gray traces represent cluster input on single dendrites. The color trace shows synapses scattered throughout the branch. (Right) IO curves obtained by scattered synaptic inputs throughout the cumulative branches on top of br1.
Figure 4: Somatic spiking response boosted by globally scattered synaptic inputs across strong sublinear dendrites. (A) Spiking activity generated by synaptic inputs under three spatial and two temporal patterns. (Left) cluster input: synapses distributed in a single dendrite. (Middle) locally scattered: synapses distributed in individual branches. (Right) globally scattered: synapses distributed across all branches. Somatic spiking activity in response to either sychronous or asynchronous 50 synaptic inputs under each spatial stimulation pattern at 50 Hz. (B) Spiking activity vs. asynchronous input under three spatial stimulation patterns with (left) 50, 100, and 150 synapses; (right) 50, 100, and 150 Hz.
@@
Given that PCs typically show a significantly high firing rate and install a large mumbler of dendrites, they receive a large number of input from Granular cells over the entire dendrite field. We then study how different spatial types of synaptic inputs can affect somatic firing through dendritic sublinearity. We selected dendrites with a sublinearity index greater than 20% as spiny dendrites and grouped them into 24 branches along the main dendrites (Figure 3A). For each branch, only one dendritic segment connects it to the parent main dendrite. To simulate clustered synaptic input, 50 synapses were distributed on a single dendrite or branch and activated synchronously. Figure 3B shows the somatic response by stimulating single dendrites (gray) and individual branches (colored) with a clustered input. Overall, the maximum membrane potential is greater when synapses are randomly distributed throughout the dendrites within a branch (scattered input) than when synapses are clustered on a single dendrite (cluster input), except for those dendrites close to the main trunk. Distance-dependent properties make proximal dendrites more responsive.
@@
To further examine how dendritic sublinear integration regulates somatic output under different spatial distributions, we considered three input strategies of synapses(Figure 4A, top), using dendritic 419 in branch 1, which has the highest peak of EPSP, as an example. When all 50 synapses are activated synchronously, spikes are induced in all distribution strategies, but locally scattered has fewer spikes (Figure 4A, middle). However, when synapses are activated asynchronously as the Poisson process, spikes are generated only under the global scattered distribution (Figure 4A, bottom). We further increased the number of synapses and stimulation frequency and found that only the global scattered distribution induced spiking activity under asynchronous activation in branch 1 (Figure 4B). In contrast, both cluster input and locally scattered input can not effectively produce spiking. However, changing the number of synapses and stimulation frequency during synchronous activation induces spike production under all distribution strategies without selectivity (Figure S3).
@@ III-C Strong sublinearity suppress bursting spike
Figure 6: Bursting complex spikes are suppressed by strong sublinearity. (A) 400 synapses cluster input in three dendrites. Somatic voltage responses triggered by each stimulation site with and without NMDARs. The triangle represents the time of stimulation. (B) Scatter plots showing the peak of EPSP and sublinearity of main dendrites with NMDA and without NMDA. Cyan scatter and dendrites indicated that somatic has a burst response. (C) Similar to B, but with 200 synapses (left) and 600 synapses (right) cluster input. (D) Strong sublinearity suppresses burst. PC installed with 1000 PF AMPA synapses stimulated at 30 Hz Poisson spikes and a CF with 200, 400 and 800 AMPA+NMDA synapses distributed on different sublinear dendrites.
@@
We then combined both PF and CF inputs to simulate a more realistic scenario. Background excitatory input was mimicked by distributing PF synapses evenly onto spiny dendrites at 30 Hz, and CF inputs were mimicked by distributing synapses onto the different sublinear regions with different numbers of synapses (Figure 6D). With 200 synapses, complex spikes occur when synapses are distributed in dendritic regions with a sublinearity of less than 20%, causing increased somatic maximum spike rate. With 400 synapses, CS can not be induced when synapses are distributed only in dendrites with sublinearity greater than 40%. When the input increases to 800 synapses, CSs occur more robustly and therefore the somatic maximum spike rate increases further. Overall, stronger sublinearity inhibits CS production. These results reveal the mechanism of bursting complex spikes highly depends on the dendritic sublinearity.
@@ III-D Dendritic sublinearity implements the feature binding
Figure 7: Sublinear dendrites implement Boolean functions. (A) 24 branches with each representing a computing unit. Synapses are binary as 0 or 1 for the truth table of AND, OR, and XOR with two inputs ($\rm X_{1}$ and $\rm X_{2}$). (B) Branches 15 & 16 for AND, and branches 1 & 3 for OR. Colored dots indicate the distribution location of synapses, blue for inactivation, and red for activation. 100 synapses with 50 Hz synchronous input on each branch. (C) Two synchronous inputs on the left, right, and both the main branches. The bar chart shows the ratio of pairs of branches for AND and OR. (D) The distribution of branches participating in AND (blue) and OR (green) with different synapse inputs. Gray areas indicate regions not participating in any Boolean operations.
@@
Figure 9: (A) Schematic diagram illustrating the sorting of dendritic branch sublinearity indices (SI) and the implementation of AND operations (cyan) with different numbers of synapses (50, 100, 200, 300) under synchronous and asynchronous input conditions. In general, it is not easy to produce AND operations when the sublinearities of x1 and x2 are both weak. Higher sublinearity indices facilitate AND operations. (B) When executing FBP and dFBP, a schematic diagram of the sublinearity of the four branches.
@@
We finally examine the functional role of dendritic sublinearity using classical computing tasks. Our PC model has 24 branches based on dendritic sublinearity, where each branch can be a single computing unit. We first consider the computing task of the Boolean function. For this, synapses are assigned with a binary value of 0 or 1. The final output is 1 if the PC fire spikes, otherwise 0 for no spike. In this context the input-output relationship can be described by a unique truth table, corresponding to a Boolean function (Figure 7A). The truth table describes three simple Boolean functions: OR, AND, and XOR. To compute this, we applied synaptic inputs on a pair of branches, went through all the pairs out of 24 branches, and recorded the somatic membrane response (Figure 7B). Examples show that the neural model can implement either the AND (branches 15 and 16), or OR (branches 1 and 3) Boolean function. However, no example of XOR was found. The particular PC has two major trunk dendrites, so we exhausted all possible pairs on both trunks and counted the percentage of pairs for each Boolean function (Figure 7C). In the left main trunk, only AND can be performed when the number of synaptic inputs is 50, and the cell starts to perform OR as the number of synapses increases. At the right main trunk, the PC implements a higher proportion of ANDs and requires a higher number of synapses to perform OR. When the two branches are from the two main trunks, the change in the number of synapses has little effect on performing AND, while more difficult to implement OR. In general, the cell prefers to perform AND, and more synaptic inputs are required to perform OR with synchronous inputs (Figure 7D). Such observations are relatively independent of the temporal input protocols and preserved when asynchronous inputs were used (Supplemental Figure S4).

### v2

@@ II-C Data analysis
We simulated the PC morphology model in NEURON 8.0. The time step of all simulation experiments was set at 0.025 ms. Data collected after simulations were saved for further analysis. In the default model, we turned off ionic currents in simulations. By removing active ion channel distributions, we aimed to simplify the model and focus on how the spatial arrangement of synapses and the inherent passive properties of dendrites shape integration, free from the influence of active currents. However, the dendritic nonlinearity is even more pronounced compared to the passive model, when active channels are turned on. To focus on the effect of nonlinearity on PC dynamics, we used the passive model as the default model.
To study dendritic nonlinearity, we record the observed EPSPs ($EPSP_{O}$), and the EPSPs produced by individual synapses are added together to obtain the expected EPSP ($EPSP_{E}$). The sublinearity index was measured as the average sublinear summation $(1-(EPSP_{O}/EPSP_{E}))*100$ with the last two points as in the dashed box in Figure 1. To investigate the temporal dependence of dendritic integration, we systematically varied the inter-spike intervals (ISI) between synaptic inputs. Then, the amplitude of all EPSPs induced by the stimuli was averaged to obtain the EPSP value elicited by the stimulus. From this, the observed values and expected values were obtained, and the sublinearity index was calculated.
@@
$F(EPSP_{E})=\frac{EPSP_{O}{max}}{1+(EPSP_{E}{50}/EPSP_{E})^{n}}+EPSP_{O}{0},$ (11)
where $n$ is the exponent factor, $EPSP_{O}{0}$ is the observed EPSP offset, and $EPSP_{O}{max}$ is the maximum observed EPSP. $EPSP_{E}{50}$ is the value of expected EPSP at which observed EPSP reaches half maximum. To characterize EPSP in detail, we introduced three measures, the peak, rise time, and width of EPSP. Peak is the maximum peak value of EPSP, and width as the distance between the points where the half-amplitude reaches before and after the peak.
To further study the effect of dendritic nonlinearity on PC firing. The dendrites with less than $20\%$ sublinearity are regarded as main dendrites, and those with more than $20\%$ are regarded as spiny dendrites, and then the spiny dendrites are divided into 24 branches, and the parent node of each branch is the main dendrite.
II-D Implementation of Feature binding problem
Feature binding problem (FBP) refers to the brain’s ability to integrate multiple attributes (e.g., color, shape, motion) into a unified percept. In the context of Purkinje cell computation, each dendritic branch is treated as an independent computational unit. Synaptic activation on a branch is represented as input as 1, while the absence of activation is represented as 0. In the simplest case, 2 out of 24 dendritic branches are chosen as inputs ($X_{1},X_{2}$). These inputs represent binary features, and the PC was tasked with performing basic Boolean operations, such as AND, OR, or XOR. This level tests the PC’s ability to compute simple logical functions based on spatially distributed inputs. Specific dendritic branches were selected to represent distinct features. In the spatial FBP task, any 4 out of 24 branches were designated as $X_{1},X_{2},X_{3},X_{4}$, where the neuron is required to bind these features into a coherent response. The neuron’s response depends not only on which branches are activated but also on the specific combination of inputs.
To incorporate temporal dynamics, we extended the FBP to a spatiotemporal setting. Here, the inputs ($X_{1},X_{2},X_{3},X_{4}$) are presented in three sequential temporal steps ($t_{1},t_{2},t_{3}$). Two features are activated in each step, and the neuron’s response depends not only on the spatial pattern of inputs but also on their timing. The neuron fires only when $X_{3}$ and $X_{4}$ are activated during the second $t_{2}$ and third $t_{3}$ steps. Any change in the temporal order or timing of activation prevents PC firing. In this way, PC is dedicated to select the spatiotemporal patterns of assigned features.
@@ III-A PC dendrites show sublinear integration
Figure 1: PC individual dendrites show sublinear integration to the somatic response. (A) (Left) 5 synapses are distributed on thin (blue) and thick (red) dendrites. (Right) Single EPSPs evoked by single-pulse synaptic stimulation; Observed EPSPs evoked with increasing synapses; Expected EPSPs as the algebraic sum of single EPSPs. (B) Sublinear relationship of the expected and observed EPSPs from (A). The dotted line represents a linear function. The sublinearity index, 49% (thin) and 15% (thick) was measured as the average sublinear summation in the dashed box. The solid line indicates fits by a Hill function. (C) The distributed sublinearity index for each dendrite. Sublinearity as a function of dendritic diameter and distance from soma.
@@
Figure 2: Temporal dependence of sublinear dendritic integration. (A) Synaptic inputs are distributed across thin (blue) and thick (red) dendritic (left). Single EPSPs were evoked using six-pulse synaptic stimulation, and the observed EPSPs were measured with an increasing number of active synapses. The expected EPSPs were calculated as the algebraic sum of the individual EPSPs (right). A pulse interval of 5 ms was used. (B) The relationship between observed and expected EPSPs at pulse intervals of 1 ms, 5 ms, and 10 ms, normalized to the maximum response. The dotted line indicates the expected linear summation, while the solid lines represent fits to the Hill function. (C) The sublinearity index is mapped across the dendritic for pulse intervals of 1 ms, 5 ms, and 10 ms, with the color scale indicating the degree of sublinearity (left). Sublinearity is plotted as a function of dendritic diameter (right).
@@
To examine whether the large dendritic depolarization could decrease the synaptic current driving force and introduce a nonlinearity that would curtail linear summation of somatic EPSPs (excitatory postsynaptic potentials) within the same PC dendrite, we studied subthreshold input-output relationships by comparing the algebraic sum of individual EPSP from five activated synapses along different yet nearby spatial locations along one dendrite, with the compound EPSP in response to simultaneous activation of multiple 2 to 5 locations (Figure 1A). The compound EPSP observed (${EPSP_{O}}$) was systematically smaller than the algebraic sum of its corresponding individual EPSP expected (${EPSP_{E}}$), especially on thin dendrites. The input-output relationship of dendritic summation using these five synapses shows sublinear trends that can be fitted by the Hill function. The sublinearity index (SI) can be calculated (the average of ${(1-(EPSP_{O}/EPSP_{E}))*100}$ using the last two points; dashed box in Figure 1B) for each dendrite. The SI of the thin dendrite (diameter $0.29\mu m$, 49%) is stronger than the thick dendrite (diameter $1.75\mu m$, 15%) as in Figure 1B. We then asked how dendritic properties regulate the SI (Figure 1C). The SI of 739 dendrites was distributed over the entire dendritic field, which shows apical dendrites exhibit stronger SIs, but trunk dendrites are weaker. SI decreases with larger diameter dendrites while increasing with the dendritic distance from the soma. Moreover, SI increased significantly with the addition of active ion channels installed on the PC model, indicating that active ion channel dynamics enhance dendritic sublinearity (Figure S1). Such observations are inherent in the observed and expected EPSPs (Figure S2). The peak, rise time, and width of EPSP are also location-dependent (Figure S3). We further observed consistent nonlinearity patterns with several additional PCs with different morphologies from various species (Figure S4), including a rat PC (NMO-00891), a mouse PC (NMO-00864), and a guinea pig PC (NMO-00610) available on public archives (www.neuromorpho.org).
@@ III-B Strong sublinearity prefers a global scatter input strategy
Figure 3: Scattered synaptic inputs on strong sublinear dendrites trigger larger somatic responses. (A) 24 branches (color-coded) with strong sublinearity are segmented in the entire morphology. Each branch groups a set of single dendrites with a sublinearity index ¿20%. The black trunk represents the main dendrites (dendritic sublinearity ¡20%). (B) Somatic EPSP in response to 50 synaptic inputs in a branch. Individual gray traces represent the response to a cluster input (50 synapses are distributed closely) on a single dendrite. The color trace shows synapses scattered throughout the branch. (C) Illustration of the somatic spiking threshold $\theta$ with varying degrees (solid lines) of dendritic sublinearity. Strong sublinearity makes soma voltage more difficult to reach the spiking threshold. (D) Input-output (IO) relationship curves change over the sum of dendritic branches. (Left) IO curves in branch 1 (br1 in (B)). Gray traces represent cluster input on single dendrites. The color trace shows synapses scattered throughout the branch. (Right) IO curves were obtained by scattered synaptic inputs throughout the cumulative branches on top of br1.
Figure 4: Somatic spiking response boosted by globally scattered synaptic inputs across strong sublinear dendrites. (A) Spiking activity generated by synaptic inputs under three spatial and two temporal patterns. (Left) cluster input: synapses distributed in a single dendrite. (Middle) locally scattered: synapses distributed in individual branches. (Right) globally scattered: synapses distributed across all branches. Somatic spiking activity in response to either synchronous or asynchronous 50 synaptic inputs under each spatial stimulation pattern at 50 Hz. (B) Spiking activity vs. asynchronous input under three spatial stimulation patterns with (left) 50, 100, and 150 synapses; (right) 50, 100, and 150 Hz. (C). Similar to B, but with synchronous input.
@@
Given that PCs typically show a significantly high firing rate and install a large mumbler of dendrites, they receive a large number of input from granule cells over the entire dendrite field. We then study how different spatial types of synaptic input can affect somatic firing through dendritic sublinearity. We selected dendrites with a sublinearity index greater than 20% as spiny dendrites and grouped them into 24 branches along the main dendrites (Figure 3A). Characteristics, the length, diameter and area, of 24 branches are consistent with the methodology used in [50], indicating significant variability among the branches and highlighting the structural diversity within the selected branches (Figure S5). For each branch, only one dendritic segment connects it to the parent main dendrite. To simulate clustered synaptic input, 50 synapses were distributed on a single dendrite or branch and activated synchronously. Figure 3B shows the somatic response by stimulating single dendrites (gray) and individual branches (colored) with a clustered input. Overall, the maximum membrane potential is greater when synapses are randomly distributed throughout the dendrites within a branch (scattered input) than when synapses are clustered on a single dendrite (cluster input), except for those dendrites close to the main trunk. Distance-dependent properties make proximal dendrites more responsive.
@@
To further examine how dendritic sublinear integration regulates somatic output under different spatial distributions, we considered three input strategies of synapses(Figure 4A, top), using dendritic 419 in branch 1, which has the highest peak of EPSP, as an example. When all 50 synapses are activated synchronously, spikes are induced in all distribution strategies, but locally scattered has fewer spikes (Figure 4A, middle). However, when synapses are activated asynchronously as the Poisson process, spikes are generated only under the global scattered distribution (Figure 4A, bottom). We further increased the number of synapses and stimulation frequency and found that only the global scattered distribution induced spiking activity under asynchronous activation in branch 1 (Figure 4B). In contrast, both cluster input and locally scattered input can not effectively produce spiking. Additionally, similar results are observed in the guinea pig (NMO_00610) PC cell model (Figure S6A). Specifically, only the globally scattered input distribution was capable of generating spikes, while both cluster input and locally scattered input failed to produce spiking activity, even at higher input rates. However, changing the number of synapses and stimulation frequency during synchronous activation induces spike production under all distribution strategies(Figure 4C).
@@ III-C Strong sublinearity suppress bursting spike
Figure 6: Bursting complex spikes are suppressed by strong sublinearity. (A) 400 synapses cluster input in three dendrites. Somatic voltage responses are triggered by each stimulation site with and without NMDARs. The triangle represents the time of stimulation. (B) Scatter plots showing the peak of EPSP and sublinearity of main dendrites with NMDA and without NMDA. Cyan scatter and dendrites indicated that somatic has a burst response. (C) Similar to B, but with 200 synapses (left) and 600 synapses (right) cluster input. (D) Strong sublinearity suppresses burst. PC installed with 1000 PF AMPA synapses stimulated at 30 Hz Poisson spikes and a CF with 200, 400, and 800 AMPA+NMDA synapses distributed on different sublinear dendrites.
@@
Purkinje cells can be categorized into two types, Zebrin positive (Z+) and Zebrin negative (Z-) [51]. Z+ PCs are characterized by lower background firing rates and weaker responses to synaptic input, while Z- PCs exhibit higher background activity and stronger input responses [52]. We combined PF and CF inputs to simulate a more realistic scenario. Background excitatory input was mimicked by distributing PF synapses evenly onto spiny dendrites at 30 Hz, and CF inputs were mimicked by distributing synapses onto the different sublinear regions with different numbers of synapses (Figure 6D). With 200 synapses, complex spikes occur when synapses are distributed in dendritic regions with a sublinearity of less than 20%, causing increased somatic maximum spike rate. With 400 synapses, CS can not be induced when synapses are distributed only in dendrites with sublinearity greater than 40%. When the input increases to 800 synapses, CSs occur more robustly and therefore the somatic maximum spike rate increases further. The similar results are observed in the guinea pig (NMO_00610) PC (Figure S6B). Overall, stronger sublinearity inhibits CS production. Under high-frequency (100 Hz) background stimulation, the occurrence of CSs is suppressed compared to low-frequency stimulation (Figure S7). Strong nonlinear responses continue to suppress the generation of CSs, regardless of the background frequency. These results reveal the mechanism of bursting complex spikes highly depends on the dendritic sublinearity.
@@ III-D Dendritic sublinearity implements the feature binding
Figure 7: Sublinear dendrites implement Boolean functions. (A) 24 branches with each representing a computing unit. Synapses are binary as 0 or 1 for the truth table of AND, OR, and XOR with two inputs ($X_{1}$ and $X_{2}$). (B) Branches 15 & 16 for AND, and branches 1 & 3 for OR. Colored dots indicate the distribution location of synapses, blue for inactivation, and red for activation. 100 synapses with 50 Hz synchronous input on each branch. (C) Two synchronous inputs on the left, right, and both the main branches. The bar chart shows the ratio of pairs of branches for AND and OR. (D) The distribution of branches participating in AND (blue) and OR (green) with different synapse inputs. Gray areas indicate regions not participating in any Boolean operations.
@@
We finally examine the functional role of dendritic sublinearity using classical computing tasks. Our PC model has 24 branches based on dendritic sublinearity, where each branch can be a single computing unit. We first consider the computing task of the Boolean function. For this, synapses are assigned with a binary value of 0 or 1. The final output is 1 if the PC fire spikes, otherwise 0 for no spike. In this context the input-output relationship can be described by a unique truth table, corresponding to a Boolean function (Figure 7A). The truth table describes three simple Boolean functions: OR, AND, and XOR. To compute this, we applied synaptic inputs on a pair of branches, went through all the pairs out of 24 branches, and recorded the somatic membrane response (Figure 7B). Examples show that the neural model can implement either the AND (branches 15 and 16), or OR (branches 1 and 3) Boolean function. However, no example of XOR was found. The particular PC has two major trunk dendrites, so we exhausted all possible pairs on both trunks and counted the percentage of pairs for each Boolean function (Figure 7C). In the left main trunk, only AND can be performed when the number of synaptic inputs is 50, and the cell starts to perform OR as the number of synapses increases. At the right main trunk, the PC implements a higher proportion of ANDs and requires a higher number of synapses to perform OR. When the two branches are from the two main trunks, the change in the number of synapses has little effect on performing AND, while more difficult to implement OR. In general, the cell prefers to perform AND, and more synaptic inputs are required to perform OR with synchronous inputs (Figure 7D). Such observations are relatively independent of the temporal input protocols and preserved when asynchronous inputs were used (Figure S8).
Figure 9: Sublinear dendrites can achieve a time-dependent feature binding. (A) The diagram illustrates three time steps ($t_{1}$, $t_{2}$, $t_{3}$) with four features distributed across dendritic branches 8, 7, 19, and 23. The feature vectors ($X_{1}$, $X_{2}$, $X_{3}$, $X_{4}$) represent activation states with 0 for inactive and 1 for active. FBP is successful if spikes occur only when the features ($X_{3}$, $X_{4}$) are activated during the time steps ($t_{2}$, $t_{3}$), indicating by True, otherwise False. Miss indicates the failure of FBP. (B) Spike count across all 216 possible combinations of feature activations, where the X, Y, and Z axes represent feature vectors from three time periods. The color scale indicates the spike count, with 5 True cases and only 1 Miss case, as shown in (A).